What is the mass of brass that requires 5200J to heat it from 23 C to 100 C? The specific heat of brass is 0.380 J/g*C.


1
Expert's answer
2021-05-24T07:19:36-0400

Solution:

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C-1 g-1)

Tf = final temperature (°C)

Ti = initial temperature (°C)

 

q = 5200 J

m = unknown

C = 0.380 J °C-1 g-1

Tf = 100°C

Ti = 23°C


Thus:

(5200 J) = m × (0.380 J °C-1 g-1) × (100 - 23)°C

(5200 J) = m × (29.26 J g-1)

m = (5200 J) / (29.26 J g-1) = 177.717 g = 177.72 g

m = 177.72 g


Answer: The mass of brass is 177.72 g